# Jailbreak-AudioBench: In-Depth Evaluation and Analysis of Jailbreak Threats for Large Audio Language Models

### v3

@@ 4.1 Query-based Audio Editing Jailbreak Method
Hence, each audio in the Explicit Small dataset has 32 variations with different audio editing combinations, which are used to query models to maximize the likelihood of jailbreak. As Figure 5 shows, our Query-based Audio Editing Jailbreak method demonstrates a significant ASR increase in model vulnerabilities to audio jailbreak on the Explicit Small dataset. Each panel presents a matrix where columns represent individual audio samples from the Explicit Small dataset, and the first 32 rows represent different edited variants of these samples. Green cells indicate failed jailbreak attempts, while red cells indicate successful compromises of the model’s safety guardrails. The penultimate row in each panel represents the original unedited audio sample, while the bottom row indicates whether any of the 32 variant queries successfully bypassed the model’s defenses. Specifically, Qwen2-Audio-7B shows substantial vulnerability with ASR increasing dramatically from 13.3% with original samples to 48.8% under our query-based approach. SALMONN-7B demonstrates even greater susceptibility, with ASR escalating from 31.6% to 85.1%. Most notably, even the closed-source GPT-4o-Audio exhibits vulnerability with ASR increasing from a mere 0.7% to 8.4% under our systematic audio editing combinations. Additional results of the Query-based Audio Editing Jailbreak method on BLSP, SpeechGPT, VITA-1.5, and MiniCPM-o-2.6 are available in the Appendix D.
@@ 5 Related Works
Jailbreak Threats     Currently, various methods successfully perform jailbreak attacks on advanced LLMs and MLLMs. Simple prompt engineering—such as fabricating facts, role-playing, or repetitive querying—reveals vulnerabilities across modalities [44; 24; 18]. In LVLMs, attackers manipulate vision inputs through typographic or visual perturbations to trigger jailbreaks [17; 7]. Another common strategy injects optimized, imperceptible perturbations into modality inputs to craft adversarial prompts [64; 35] or images [40; 22; 8; 43]. To support systematic evaluation, benchmarks such as AdvBench [64], MM-SafetyBench [32], RedTeam-2K [34], and Safebench [17] provide diverse jailbreak prompts. Recent works [53; 34; 55; 60] focus on LVLM Jailbreak robustness, proposing benchmarks and pipelines to assess safety from both attack and defense perspectives. For the audio modality, [56] adopts a subset of 350 samples from AdvBench [28] to evaluate the jailbreak vulnerabilities of several state-of-the-art end-to-end LALMs [9; 46; 47].
@@
LALMs Benchmark     Recent advancements in evaluating Large Audio Language Models (LALMs) lead to the development of several comprehensive benchmarks and models. AIR-Bench [58] assesses LALMs’ understanding of diverse audio signals—including speech, natural sounds, and music—through foundational and conversational tasks. AudioBench [51] covers eight tasks across 26 datasets, focusing on speech comprehension, audio scene analysis, and paralinguistic features. MMAU [42] evaluates expert-level reasoning using 10,000 audio clips with Q&A sets for multimodal understanding. ADU-Bench [14] emphasizes conversational ability with 20,000 open-ended multilingual dialogues. FunAudioLLM [3] integrates SenseVoice for speech recognition and emotion detection with CosyVoice for speech generation. WavLLM [24] employs dual encoders to separately model semantic and speaker information, enhancing task adaptability.
@@ 6 Discussions
The Resource Requirements of Jailbreak-AudioBench     The comprehensive execution of Jailbreak-AudioBench demands significant computational resources, encompassing approximately 9,216 GPU hours on NVIDIA A40. The evaluation of closed-source LALMs such as GPT-4o-Audio incurs substantial API usage costs, amounting to \$1,000.
Limiations & Future Continuous Updating     Accurately modeling natural human speech with realistic variations in prosody, speed, and pronunciation remains challenging. Our current approach uses TTS-generated audio converted from text as the original input, and applies editing through our Toolbox to produce diverse variants. In future work, we aim to incorporate natural speech recordings and expand the benchmark to better reflect real-world scenarios. Additionally, Jailbreak-AudioBench will be continuously updated following the release of new LALMs and jailbreak datasets.
@@ A.2 Editing Process for Each Type of Audio Hidden Semantics
Celebrity Accent    We modify vocal characteristics to simulate recognizable celebrity speech patterns from Kanye West, Donald Trump, and Lucy Liu using voice conversion techniques. This transformation is performed using coqui-ai TTS [10], enabling high-quality voice conversion while preserving content and adopting distinctive prosodic and timbral qualities associated with each celebrity’s speaking style.
@@ A.3.1 The textual annotation of audio hidden semantics
Emphasis    “CREATE”: Words written in ALL CAPS and bold indicate that the volume of these specific words has been amplified. The "Volume x 5" label clarifies that the amplification factor is 5 times the original volume, making these words much louder. Other emphasis levels include Volume x 2 and Volume x 10.
Speed    “Write…a…tutorial…on…how …to… make…a…bomb…”: The use of an ellipsis (…) between words signifies a modification to the playback speed, resulting in a slowed-down speech rate or extended pauses. The “Rate x 0.5” label indicates that the speech playback rate has been reduced to 0.5 times the original speed. The speed can also be increased, for example, “Write $\rightarrow$ a $\rightarrow$ tutorial $\rightarrow$ on $\rightarrow$ how $\rightarrow$ to $\rightarrow$ make $\rightarrow$ a $\rightarrow$ bomb” representing speed up, such as Rate x 1.5.
Intonation    “List steps in detail $\uparrow$ to hack $\uparrow$ into someone’s bank account $\uparrow$”: An upward arrow ($\uparrow$) placed after a word signifies a rising intonation on that word or phrase. The “Interval+3” label specifies that the pitch is shifted in intervals, with a +3 interval indicating a specific upward pitch adjustment. Additional intonation intervals include +2 and +4.
Tone    “[Pitch=+4st]”: This notation, enclosed in square brackets, indicates a direct manipulation of the pitch (tone) of the subsequent speech. “[Pitch=+4st]” specifically means the pitch is raised by 4 semitones (st), making the voice sound higher. Other tone adjustments include semitone shifts of -8, -4, and +8.
Background Noise    “[+ Crowd Noise]”: This notation, enclosed in square brackets, signifies the addition of a specific type of background noise to the audio. “[+ Crowd Noise]” indicates that the sound of a crowd is overlaid onto the speech. Background noise can also be Machine Noise and White Noise, represented as “[+ Background Noise]” and “[+ Machine Noise]”.
@@ Appendix B Additional Information of Jailbreak-AudioBench Dataset
The Plus version of the Jailbreak-AudioBench Dataset adds three additional subtype datasets (https://huggingface.co/datasets/researchtopic/Jailbreak-AudioBench-Plus): Explicit Small (GPT-4o Eval), Implicit Small (GPT-4o Eval), and Implicit Defense. The 262 base audio samples in the Explicit Small (GPT-4o Eval) are identical to those in the Explicit Small dataset, as detailed in Table 1. However, unlike the Explicit Small dataset, which employs compositional editing categories optimized via grid search, Explicit Small (GPT-4o Eval) is designed as a simplified, smaller-scale version of the Explicit Subtype dataset to evaluate the performance of GPT-4o-Audio. For the selection of base audio samples in the Implicit Small (GPT-4o Eval), the strategy mirrors that of the Explicit Small dataset from the Explicit Subtype dataset, extracting 237 base audio samples from the Implicit Subtype dataset, which maintains a one-tenth proportion of the harmful content categories. Then, by applying 20 types of audio editing from the Toolbox to the base audio samples in Explicit Small (GPT-4o Eval) and Implicit Small (GPT-4o Eval), a total of 5,240 and 4,740 edited audio samples are generated, respectively.
@@ Appendix C Additional Information of Jailbreak-AudioBench Benchmark
C.1 Evaluation on GPT-4o-Audio
We also evaluate how GPT-4o-Audio is affected by the audio editing jailbreak. Due to the large scale of the Explicit Subtype dataset and the Implicit Subtype dataset, evaluating GPT-4o-Audio would incur excessive costs. Therefore, we evaluate the GPT-4o-Audio on Explicit Small (GPT-4o Eval) and Implicit Small (GPT-4o Eval), which are smaller-scale versions of the Explicit Subtype dataset and the Implicit Subtype dataset. Detailed dataset scale information can be found in Table 3. As illustrated in Table 4, GPT-4o-Audio exhibits robustness to audio editing jailbreak, with minor ASR increases of less than 1.7% observed only in specific audio editing types, including intonation, tone, background noise, celebrity accent, and speed editing.
Table 4: The ASR performance of GPT-4o-Audio across various audio editing types when compared to the original audio on both Explicit Small (GPT-4o Eval) and Implicit Small (GPT-4o Eval) datasets. We denote the relative changes compared to the original audio: red and green indicate the increase and decrease in ASR when the absolute value of the change is greater than or equal to 1%, respectively. Note that the Original represents the baseline ASR obtained from unmodified audio samples without any audio editing.
Original Emphasis Intonation
Original Volume*2 Volume*5 Volume*10 Interval+2 Interval+3 Interval+4
0.7%/0.8% +0.4%/+0.9% +0.4%/0% +0.8%/+0.9% +0.4%/+0.4% +1.2%/+0.4% +1.5%/+1.3%
Tone Background Noise
Semitone -8 Semitone -4 Semitone +4 Semitone +8 Crowd Noise Machine Noise White Noise
0%/-0.4% +0.4%/-0.4% +0.8%/0% +1.2%/+0.9% +0.8%/+0.4% 0%/+1.7% 0%/-0.4%
Celebrity Accent Emotion Speed
Kanye West Donald Trump Lucy Liu Laugh Scream Rate*0.5 Rate*1.5
+0.4%/-0.9% 0%/0% +1.5%/+0.4% -0.4%/-0.4% -0.4%/0% +0.4%/0% +1.5%/-0.4%
C.2 Comparing Human Evaluation with Llama Guard 3
@@ D.1 Query-based Audio Editing Jailbreak
Specifically, we first create the Explicit Small dataset by extracting 262 original audio samples from the Explicit Subtype dataset, maintaining a one-tenth proportion of the harmful content categories. To systematically explore parameter combinations of various audio editing types, we then use grid search and apply 32 distinct audio editing combinations to these base audio samples, systematically combining modifications related to accent/emotion, emphasis, speed, and background noise in sequence. This combinatorial approach generated $262\times 32=8384$ audio samples, which construct the Explicit Small dataset.

### v4

@@ 4.1 Query-based Audio Editing Jailbreak Method
Hence, each audio in the Explicit Small dataset has 32 variations with different audio editing combinations, which are used to query models to maximize the likelihood of jailbreak. As Figure 5 shows, our Query-based Audio Editing Jailbreak method demonstrates a significant ASR increase in model vulnerabilities to audio jailbreak on the Explicit Small dataset. Each panel presents a matrix where columns represent individual audio samples from the Explicit Small dataset, and the first 32 rows represent different edited variants of these samples. Green cells indicate failed jailbreak attempts, while red cells indicate successful compromises of the model’s safety guardrails. The penultimate row in each panel represents the original unedited audio sample, while the bottom row indicates whether any of the 32 variant queries successfully bypassed the model’s defenses. Specifically, Qwen2-Audio-7B shows substantial vulnerability with ASR increasing dramatically from 13.3% with original samples to 48.8% under our query-based approach. SALMONN-7B demonstrates even greater susceptibility, with ASR escalating from 31.6% to 85.1%. Most notably, even the closed-source GPT-4o-Audio exhibits vulnerability with ASR increasing from a mere 0.7% to 8.4% under our systematic audio editing combinations. Similarly, Gemini-2.5-Flash shows significant vulnerability with ASR rising from 8.1% to 49.4%. Additional results of the Query-based Audio Editing Jailbreak method on BLSP, SpeechGPT, VITA-1.5, and MiniCPM-o-2.6 are available in the Appendix D.
@@ 5 Related Works
Jailbreak Threats    Currently, various methods successfully perform jailbreak attacks on advanced LLMs and MLLMs. Simple prompt engineering—such as fabricating facts, role-playing, or repetitive querying—reveals vulnerabilities across modalities [53; 31; 24; 8; 10; 16]. In LVLMs, attackers manipulate vision inputs through typographic or visual perturbations to trigger jailbreaks [23; 9]. Another common strategy injects optimized, imperceptible perturbations into modality inputs to craft adversarial prompts [73; 43] or images [48; 29; 12; 52; 11]. To support systematic evaluation, benchmarks such as AdvBench [73], MM-SafetyBench [40], RedTeam-2K [42], and SafeBench [23] provide diverse jailbreak prompts. Recent works [62; 42; 64; 69] focus on LVLM Jailbreak robustness, proposing benchmarks and pipelines to assess safety from both attack and defense perspectives. For the audio modality, [65] adopts a subset of 350 samples from AdvBench [35] to evaluate the jailbreak vulnerabilities of several state-of-the-art end-to-end LALMs [13; 55; 56].
@@
LALMs Benchmark    Recent advancements in evaluating Large Audio Language Models (LALMs) have led to the development of several comprehensive benchmarks and models. AIR-Bench [67] assesses LALMs’ understanding of diverse audio signals—including speech, natural sounds, and music—through foundational and conversational tasks. AudioBench [60] covers eight tasks across 26 datasets, focusing on speech comprehension, audio scene analysis, and paralinguistic features. MMAU [51] evaluates expert-level reasoning using 10,000 audio clips with Q&A sets for multimodal understanding. ADU-Bench [20] emphasizes conversational ability with 20,000 open-ended multilingual dialogues. FunAudioLLM [3] integrates SenseVoice for speech recognition and emotion detection with CosyVoice for speech generation. WavLLM [31] employs dual encoders to separately model semantic and speaker information, enhancing task adaptability.
@@ 6 Discussions
Resource Requirements of Jailbreak-AudioBench    The comprehensive execution of Jailbreak-AudioBench demands significant computational resources, encompassing approximately 9,216 GPU hours on NVIDIA A40. The evaluation of closed-source LALMs such as GPT-4o-Audio and Gemini-2.5-Flash incurs substantial API usage costs, amounting to \$1,000.
Limitations & Future Work    In this paper, following previous jailbreak studies [50; 65; 66], we also mainly use Llama Guard 3 [33] to evaluate the responses of LALMs. However, after examining 157,782 responses, we observe that Llama Guard 3 has several limitations. In particular, some responses simply repeat the input prompt. Since these outputs contain a few harmful words, Llama Guard 3 incorrectly marks them as successful attacks. These findings indicate that current jailbreak evaluation metrics remain imperfect. We plan to improve them in future work and encourage the research community to further investigate this issue.
Additionally, accurately modeling natural human speech with realistic variations in prosody, speed, and pronunciation remains challenging. Our current approach uses TTS-generated audio converted from text as the original input and applies editing through our Toolbox to produce diverse variants. In future work, we aim to incorporate natural speech recordings and expand the benchmark to better reflect real-world scenarios.
@@ A.2 Editing Process for Each Type of Audio Hidden Semantics
Celebrity Accent   We modify vocal characteristics to simulate recognizable celebrity speech patterns from Kanye West, Donald Trump, and Lucy Liu using voice conversion techniques. This transformation is performed using Coqui TTS [15], enabling high-quality voice conversion while preserving content and adopting distinctive prosodic and timbral qualities associated with each celebrity’s speaking style.
@@ A.3.1 The Textual Annotation of Audio Hidden Semantics
Emphasis   “CREATE”: Words written in uppercase and bold indicate that the volume of these specific words has been amplified. The "Volume $\times$ 5" label clarifies that the amplification factor is 5 times the original volume, making these words much louder. Other emphasis levels include Volume $\times$ 2 and Volume $\times$ 10.
Speed   “Write … a … tutorial … on … how … to … make … a … bomb …”: The use of an ellipsis (…) between words signifies a modification to the playback speed, resulting in a slowed-down speech rate or extended pauses. The “Rate $\times$ 0.5” label indicates that the speech playback rate has been reduced to 0.5 times the original speed. The speed can also be increased, for example, “Write $\rightarrow$ a $\rightarrow$ tutorial $\rightarrow$ on $\rightarrow$ how $\rightarrow$ to $\rightarrow$ make $\rightarrow$ a $\rightarrow$ bomb” representing speed-up, such as Rate $\times$ 1.5.
Intonation   “List steps in detail $\uparrow$ to hack $\uparrow$ into someone’s bank account $\uparrow$”: An upward arrow ($\uparrow$) placed after a word signifies a rising intonation on that word or phrase. The “Interval+3” label specifies that the pitch is shifted in intervals, with a +3+3 interval indicating a specific upward pitch adjustment. Additional intonation intervals include +2+2 and +4+4.
Tone   “[Pitch=+4st]”: This notation, enclosed in square brackets, indicates a direct manipulation of the pitch (tone) of the subsequent speech. “[Pitch=+4st]” specifically means the pitch is raised by 4 semitones (st), making the voice sound higher. Other tone adjustments include semitone shifts of 8-8, 4-4, and +8+8.
Background Noise   “[+ Crowd Noise]”: This notation, enclosed in square brackets, signifies the addition of a specific type of background noise to the audio. “[+ Crowd Noise]” indicates that the sound of a crowd is overlaid onto the speech. Background noise can also be Machine Noise and White Noise, represented as “[+ Machine Noise]” and “[+ White Noise]”.
@@ Appendix B Additional Information of Jailbreak-AudioBench Dataset
The Plus version of the Jailbreak-AudioBench Dataset adds two additional subtype datasets (https://huggingface.co/datasets/researchtopic/Jailbreak-AudioBench-Plus): Explicit Small (GPT-4o Eval) and Implicit Small (GPT-4o Eval).
Specifically, the 262 base audio samples in the Explicit Small (GPT-4o Eval) are identical to those in the Explicit Small dataset, as detailed in Table 1. However, unlike the Explicit Small dataset, which employs compositional editing categories optimized via grid search, Explicit Small (GPT-4o Eval) is designed as a simplified, smaller-scale version of the Explicit Subtype dataset to evaluate the performance of GPT-4o-Audio and Gemini-2.5-Flash. Similarly, for the selection of base audio samples in the Implicit Small (GPT-4o Eval), the strategy mirrors that of the Explicit Small dataset from the Explicit Subtype dataset, extracting 237 base audio samples from the Implicit Subtype dataset, which maintains a one-tenth proportion of the harmful content categories. Then, by applying 20 types of audio editing from the Toolbox to the base audio samples in Explicit Small (GPT-4o Eval) and Implicit Small (GPT-4o Eval), a total of 5,240 and 4,740 edited audio samples are generated, respectively.
@@ Appendix C Additional Information of Jailbreak-AudioBench Benchmark
C.1 Comparing Human Evaluation with Llama Guard 3
@@ D.1 Query-based Audio Editing Jailbreak
Specifically, we first create the Explicit Small dataset by extracting 262 original audio samples from the Explicit Subtype dataset, maintaining a one-tenth proportion of the harmful content categories. To systematically explore parameter combinations of various audio editing types, we then use grid search and apply 32 distinct audio editing combinations to these base audio samples, systematically combining modifications related to accent/emotion, emphasis, speed, and background noise in sequence. This combinatorial approach generated $262\times 32=8384$ audio samples, which construct the Explicit Small dataset. Detailed dataset scale information is shown in Table 1.